ID |
Date |
Author |
Target Material |
Target Oven W.O# |
Source |
Status |
Subject |
Text |
 |
529
|
Saturday, February 13, 2021, 08:14 |
Darwin Ortiz Rosales | SiC | 41992-1 | TARGET BOAT | Done | Sintering of SiC discs for SEM | Saturday, February 13, 2021, 08:10
Ta boat loaded with 80 discs |
|
528
|
Thursday, February 11, 2021, 10:57 |
Marla Cervantes | | NA | | Done | Re coating Test 1 | Thursday, February 10, 2021, 12:00
Objective: Heat a Re wire to evaporate
it and deposit a layer of Re on the inner |
|
527
|
Tuesday, February 09, 2021, 06:48 |
Aaron Schmidt | | NA | | Done | Ta coating test | Monday, February 08, 2021, 17:00: Installed
brackets w/ Ta wire and a small piece of
Re to test voltage drop. |
|
526
|
Monday, February 01, 2021, 13:57 |
Darwin Ortiz Rosales | SiC | 40282-4 | TARGET BOAT | Done | Sintering of SiC discs for SEM | Monday, February 01, 2021, 13:52
|
|
525
|
Tuesday, November 24, 2020, 12:56 |
Marla Cervantes | | 44139-8-SK | TARGET BOAT | Done | Graphite casting conditioning | Tuesday, November 24, 2020, 09:23
A batch of 60 discs of 18 mm diameter
were loaded in a tantalum tube. The tantalum |
   |
524
|
Tuesday, November 24, 2020, 08:13 |
Aaron Schmidt | | | | | Flex - line HS installed | Monday, November 23, 2020, 17:01: Flex-line
HS installed.
Tuesday, November 24, 2020, |
|
523
|
Thursday, November 05, 2020, 15:13 |
Ferran Boix Pamies | Empty | 41882-2 | TARGET BOAT | Done | cleaning treatment. | 2020-11-05 15:44 empty and naked
boat in for heating up system and clean up.
Turbo on. |
|
522
|
Wednesday, November 04, 2020, 17:31 |
Ferran Boix Pamies | Empty | 41882-2 | TARGET BOAT | Done | Parasitic target graphite outgassing. | 2020-11-04 17:31: target installed
with boat containing graphite container for
outgassing. Roughing pump on. |
|
521
|
Monday, November 02, 2020, 18:32 |
Ferran Boix Pamies | Empty | 44139-7-SK | TARGET BOAT | Done | HEA heat treatment | 2020-11-02 6:32: roughing on
2020-11-02 7:21: Turbo on for a while,
IGP1:1.8e-5T . water leak check ok. |
 |
520
|
Tuesday, October 27, 2020, 10:25 |
Ferran Boix Pamies | Empty | 44139-7-SK | TARGET BOAT | Done | parasitic target commissioning | 2020-10-27 10:24 setup introduced
yesterday. Pumping over night IGP1: 5.7e-7T.
Water leak check ok. Performing heating elements |
|
519
|
Wednesday, October 21, 2020, 14:39 |
Aaron Schmidt | Nb | 44508-2 | LP-SIS | In progress | Conditioning Nb#9 | Nb#9 LP-SIS target description:
|
|
518
|
Friday, October 16, 2020, 16:31 |
Luca Egoriti | Empty | | TARGET BOAT | In progress | HPOG annealing | Friday, October 16, 2020, 11:30
Target mounted in EVAP2. Pump
down started |
|
517
|
Thursday, October 15, 2020, 18:51 |
Ferran Boix Pamies | Empty | 44139-7-SK | TARGET BOAT | Done | heat gun plus parasitic target thermal comissioning. | 2020-10-15 18:51 : Setup installed
during the day. Pumped down to IGP1:1.5e-5T
, water leak check ok. Ramping up to 100A. |
|
516
|
Tuesday, October 13, 2020, 11:15 |
Ferran Boix Pamies | Empty | 44139-7-SK | TARGET BOAT | Done | foil heating | 2020-10-13 11:13 Target was left
under vacuum over the weekend. IGP1: 1.6e-7T
water on- leak check ok. Ramping up 2A/30s |
|
515
|
Friday, October 09, 2020, 09:50 |
Ferran Boix Pamies | Empty | 44139-7-SK | TARGET BOAT | Done | Tests, radiation gun for parasitic target | 2020-10-09 9:48 The target was installed
the day before and left pumping over night.
IGP1=4.7e-7Torr. leak check when water on, |
|
514
|
Wednesday, October 07, 2020, 12:25 |
Aaron Schmidt | | | | In progress | IEVAP2:TP1 Swap out. | Wednesday, October 07, 2020
TP1 was not spinning up past 4krpm.
Stalling & tripping out. Examined with |
|
513
|
Tuesday, October 06, 2020, 14:23 |
Aaron Schmidt | Ta | 44508-1 | LP-SIS | In progress | Conditioning Ta#61 | Ta#61 LP-SIS target description:
Total Ta foils = 470 [6x20 packs |
|
512
|
Monday, September 28, 2020, 11:53 |
John Wong | np-SiC | | TARGET BOAT | In progress | np-SiC carbonization | Monday, September 28, 2020, 11:49
6 discs of np-SiC were loaded to
the target boat wrapped with 3 layers of |
|
511
|
Monday, September 28, 2020, 09:36 |
John Wong | Nb | | TARGET BOAT | Done | Heat Treatment of Nb for Edward |
|
|
510
|
Friday, September 25, 2020, 10:32 |
Darwin Ortiz Rosales | Nb | | TARGET BOAT | Done | Heat Treatment of Nb for Edward | Friday, September 25, 2020, 10:30
Target boat with a layer of |
|